OpenStudy (anonymous):

Calculate lnx-ln(x-3)=ln4

OpenStudy (anonymous):

I managed to get x=-10, is this right?

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Using rules of logs I divided lnx-ln(x-3) which gave ln(x/x-3)=ln4 I then cancelled out the lns and factorised the x giving \[x\frac{ 1 }{ 1-3 }=4\] Then divided 4 by -0.4, I have a feeling this is wrong though

OpenStudy (anonymous):